OpenStudy (anonymous):

I choose C am I right If g(x) = 2(x - 4), find the value of x if g(x) = 20. A) 32 B) 12 C) 14 D) 10

OpenStudy (azureilai):

yes

OpenStudy (anonymous):



OpenStudy (anonymous):

This is how I solved it is it right?

OpenStudy (azureilai):

That is one way to solve it. How I did it was 20=2(x-4) 10=x-4 x=14 I just divided the 2 first instead of distributing it. But you arrive at the same answer.

OpenStudy (anonymous):

Okay thanks